# Imaging coupled vibrational, rotational, and electronic wave packet dynamics in a triatomic molecule

### v1

@@ Acknowledgements.
We thank Charles Fehrenbach for taking care of the PULSAR laser operation. We are grateful to the technical staff of the J.R. Macdonald Laboratory for their support. This work and the operation of the J.R. Macdonald Laboratory are supported by the Chemical Sciences, Geosciences, and Biosciences Division, Office of Basic Energy Sciences, Office of Science, U.S. Department of Energy, Grant no. DE-FG02-86ER13491. The PULSAR laser was provided by Grant no. DE-FG02-09ER16115, S.B. is supported by Grant no. DE-SC0020276 from the same funding agency, A.S.V. by the National Science Foundation Grant no. PHYS-1753324. S.J. is funded by the Deutscher Akademischer Austauschdienst (DAAD) program for her summer research at Kansas State University.References
- Zewail [2000] A. H. Zewail, Femtochemistry: Atomic-Scale Dynamics of the Chemical Bond, The Journal of Physical Chemistry A 104, 5660 (2000), publisher: American Chemical Society.
@@
- Boutet and Williams [2010] S. Boutet and G. J. Williams, The coherent x-ray imaging (cxi) instrument at the linac coherent light source (lcls), New Journal of Physics 12, 035024 (2010).
@@
- Ware et al. [2019] M. R. Ware, J. M. Glownia, N. Al-Sayyad, J. T. O’Neal, and P. H. Bucksbaum, Characterizing dissociative motion in time-resolved x-ray scattering from gas-phase diatomic molecules, Phys. Rev. A 100, 033413 (2019).
@@
- Guo et al. [2024] Z. Guo, Z. Zhang, Y. Deng, J. Wang, D. Ye, J. Liu, and Y. Liu, Probing double ionization with bicircular laser fields, Phys. Rev. Lett. 132, 143201 (2024).
@@
- Rudenko et al. [2006] A. Rudenko, T. Ergler, B. Feuerstein, K. Zrost, C. D. Schröter, R. Moshammer, and J. Ullrich, Real-time observation of vibrational revival in the fastest molecular system, Chemical Physics Electron Correlation and Multimode Dynamics in Molecules, 329, 193 (2006).
@@
- Felker and Zewail [1985b] P. M. Felker and A. H. Zewail, Dynamics of intramolecular vibrational‐energy redistribution (IVR). II. Excess energy dependence, The Journal of Chemical Physics 82, 2975 (1985b).
- Felker and Zewail [1985c] P. M. Felker and A. H. Zewail, Dynamics of intramolecular vibrational‐energy redistribution (IVR). III. Role of molecular rotations, The Journal of Chemical Physics 82, 2994 (1985c).
- Felker et al. [1985] P. M. Felker, W. R. Lambert, and A. H. Zewail, Dynamics of intramolecular vibrational‐energy redistribution (IVR). IV. Excess energy dependence, t‐stilbene, The Journal of Chemical Physics 82, 3003 (1985).
@@
- Karashima et al. [2021] S. Karashima, Y.-I. Suzuki, and T. Suzuki, Ultrafast Extreme Ultraviolet Photoelectron Spectroscopy of Nonadiabatic Photodissociation of CS2 from 1B2 (1u+) State: Product Formation via an Intermediate Electronic State, J. Phys. Chem. Lett. 12, 3755 (2021).
@@
- Yang et al. [2016] J. Yang, M. Guehr, X. Shen, R. Li, T. Vecchione, R. Coffee, J. Corbett, A. Fry, N. Hartmann, C. Hast, K. Hegazy, K. Jobe, I. Makasyuk, J. Robinson, M. S. Robinson, S. Vetter, S. Weathersby, C. Yoneda, X. Wang, and M. Centurion, Diffractive imaging of coherent nuclear motion in isolated molecules, Phys. Rev. Lett. 117, 153002 (2016).
@@
- Lévêque et al. [2013] C. Lévêque, A. Komainda, R. Taïeb, and H. Köppel, Ab initio quantum study of the photodynamics and absorption spectrum for the coupled 11A2 and 11B1 states of SO2, The Journal of Chemical Physics 138, 044320 (2013).
@@
- Chen et al. [2023] L. Chen, E. Wang, W. Zhao, M. Gong, X. Shan, and X. Chen, Fragmentation of SO2q+ (q = 2–4) induced by 1 keV electron collision, The Journal of Chemical Physics 158, 054301 (2023).
@@
- Rösch et al. [2023] D. Rösch, Y. Xu, H. Guo, X. Hu, and D. L. Osborn, So2 photodissociation at 193 nm directly forms s(3p) + o2(3g–): Implications for the archean atmosphere on earth, The Journal of Physical Chemistry Letters 14, 3084 (2023).
@@
- Matsuda et al. [2014] A. Matsuda, E. J. Takahashi, and A. Hishikawa, Time-resolved laser Coulomb explosion imaging using few-cycle intense laser pulses: Application to exploding CS2 in highly charged states, Journal of Electron Spectroscopy and Related Phenomena 195, 327 (2014).
@@
- Makhija et al. [2016] V. Makhija, X. Ren, D. Gockel, A.-T. Le, and V. Kumarappan, Orientation Resolution through Rotational Coherence Spectroscopy (2016), arXiv:1611.06476 [physics].
- Wangjam et al. [2021] T. N. Wangjam, H. V. S. Lam, and V. Kumarappan, Strong-field ionization of the triplet ground state of , Phys. Rev. A 104, 043112 (2021).
@@
- Jahnke et al. [2024] T. Jahnke, S. Mai, S. Bhattacharyya, K. Chen, R. Boll, M. E. Castellani, S. Dold, A. Duley, U. Frühling, A. E. Green, M. Ilchen, R. Ingle, G. Kastirke, H. V. S. Lam, F. Lever, D. Mayer, T. Mazza, T. Mullins, Y. Ovcharenko, B. Senfftleben, F. Trinter, A. T. Noor, S. Usenko, A. S. Venkatachalam, A. Rudenko, D. Rolles, M. Meyer, H. Ibrahim, and M. Gühr, X-ray coulomb explosion imaging reveals role of molecular structure in internal conversion (2024), arXiv:2405.15367 [physics.chem-ph] .

### v3

@@ Acknowledgements.
We thank Charles Fehrenbach for taking care of PULSAR laser operation. We are grateful to the technical staff of the J.R. Macdonald Laboratory for their support. This work and the operation of the J.R. Macdonald Laboratory are supported by the Chemical Sciences, Geosciences, and Biosciences Division, Office of Basic Energy Sciences, Office of Science, U.S. Department of Energy, Grant no. DE-FG02-86ER13491. The PULSAR laser was provided by Grant no. DE-FG02-09ER16115, S.B. is supported by Grant no. DE-SC0020276 from the same funding agency, A.S.V. by the National Science Foundation Grant no. PHYS-2409365. S.J. was funded by the RISE Worldwide program of German Academic Exchange Service (DAAD) for her summer research at Kansas State University.References
- Zewail [2000] A. H. Zewail, Femtochemistry: Atomic-Scale Dynamics of the Chemical Bond, The Journal of Physical Chemistry A 104, 5660 (2000).
@@
- Minitti et al. [2015] M. P. Minitti, J. M. Budarz, A. Kirrander, J. S. Robinson, D. Ratner, T. J. Lane, D. Zhu, J. M. Glownia, M. Kozina, H. T. Lemke, M. Sikorski, Y. Feng, S. Nelson, K. Saita, B. Stankus, T. Northey, J. B. Hastings, and P. M. Weber, Imaging Molecular Motion: Femtosecond X-Ray Scattering of an Electrocyclic Chemical Reaction, Phys. Rev. Lett. 114, 255501 (2015).
- Ruddock et al. [2019] J. M. Ruddock, H. Yong, B. Stankus, W. Du, N. Goff, Y. Chang, A. Odate, A. M. Carrascosa, D. Bellshaw, N. Zotev, M. Liang, S. Carbajo, J. Koglin, J. S. Robinson, S. Boutet, A. Kirrander, M. P. Minitti, and P. M. Weber, A deep UV trigger for ground-state ring-opening dynamics of 1,3-cyclohexadiene, Science Advances 5, eaax6625 (2019).
- Liu et al. [2020] Y. Liu, S. L. Horton, J. Yang, J. P. F. Nunes, X. Shen, T. J. A. Wolf, R. Forbes, C. Cheng, B. Moore, M. Centurion, K. Hegazy, R. Li, M.-F. Lin, A. Stolow, P. Hockett, T. Rozgonyi, P. Marquetand, X. Wang, and T. Weinacht, Spectroscopic and structural probing of excited-state molecular dynamics with time-resolved photoelectron spectroscopy and ultrafast electron diffraction, Phys. Rev. X 10, 021016 (2020).
- Polli et al. [2010] D. Polli, P. Altoè, O. Weingart, K. M. Spillane, C. Manzoni, D. Brida, G. Tomasello, G. Orlandi, P. Kukura, R. A. Mathies, M. Garavelli, and G. Cerullo, Conical intersection dynamics of the primary photoisomerization event in vision, Nature 467, 440 (2010).
- Wörner et al. [2011] H. J. Wörner, J. B. Bertrand, B. Fabre, J. Higuet, H. Ruf, A. Dubrouil, S. Patchkovskii, M. Spanner, Y. Mairesse, V. Blanchet, E. Mével, E. Constant, P. B. Corkum, and D. M. Villeneuve, Conical Intersection Dynamics in NO2 Probed by Homodyne High-Harmonic Spectroscopy, Science 334, 208 (2011).
- Levine and Martínez [2007] B. G. Levine and T. J. Martínez, Isomerization Through Conical Intersections, Annual Review of Physical Chemistry 58, 613 (2007).
@@
- Rudenko et al. [2006] A. Rudenko, T. Ergler, B. Feuerstein, K. Zrost, C. D. Schröter, R. Moshammer, and J. Ullrich, Real-time observation of vibrational revival in the fastest molecular system, Chemical Physics 329, 193 (2006).
@@
- Rosca-Pruna and Vrakking [2001] F. Rosca-Pruna and M. J. J. Vrakking, Experimental observation of revival structures in picosecond laser-induced alignment of , Phys. Rev. Lett. 87, 153902 (2001).
- Lee et al. [2006] K. F. Lee, D. M. Villeneuve, P. B. Corkum, A. Stolow, and J. G. Underwood, Field-free three-dimensional alignment of polyatomic molecules, Phys. Rev. Lett. 97, 173001 (2006).
- Yang et al. [2016a] J. Yang, M. Guehr, T. Vecchione, M. S. Robinson, R. Li, N. Hartmann, X. Shen, R. Coffee, J. Corbett, A. Fry, K. Gaffney, T. Gorkhover, C. Hast, K. Jobe, I. Makasyuk, A. Reid, J. Robinson, S. Vetter, F. Wang, S. Weathersby, C. Yoneda, M. Centurion, and X. Wang, Diffractive imaging of a rotational wavepacket in nitrogen molecules with femtosecond megaelectronvolt electron pulses, Nature Communications 7, 11232 (2016a).
- Fang and Gibson [2007] L. Fang and G. N. Gibson, Investigating excited electronic states of and produced by strong-field ionization using vibrational wave packets, Phys. Rev. A 75, 063410 (2007).
- Yang et al. [2016b] J. Yang, M. Guehr, X. Shen, R. Li, T. Vecchione, R. Coffee, J. Corbett, A. Fry, N. Hartmann, C. Hast, K. Hegazy, K. Jobe, I. Makasyuk, J. Robinson, M. S. Robinson, S. Vetter, S. Weathersby, C. Yoneda, X. Wang, and M. Centurion, Diffractive imaging of coherent nuclear motion in isolated molecules, Phys. Rev. Lett. 117, 153002 (2016b).
@@
- Felker and Zewail [1985b] P. M. Felker and A. H. Zewail, Dynamics of intramolecular vibrational‐energy redistribution (IVR). III. Role of molecular rotations, The Journal of Chemical Physics 82, 2994 (1985b).
@@
- Karashima et al. [2021] S. Karashima, Y.-I. Suzuki, and T. Suzuki, Ultrafast Extreme Ultraviolet Photoelectron Spectroscopy of Nonadiabatic Photodissociation of CS2 from 1B2() State: Product Formation via an Intermediate Electronic State, J. Phys. Chem. Lett. 12, 3755 (2021).
@@
- Bocharova et al. [2011] I. A. Bocharova, A. S. Alnaser, U. Thumm, T. Niederhausen, D. Ray, C. L. Cocke, and I. V. Litvinyuk, Time-resolved coulomb-explosion imaging of nuclear wave-packet dynamics induced in diatomic molecules by intense few-cycle laser pulses, Phys. Rev. A 83, 013417 (2011).
- De et al. [2010] S. De, I. A. Bocharova, M. Magrakvelidze, D. Ray, W. Cao, B. Bergues, U. Thumm, M. F. Kling, I. V. Litvinyuk, and C. L. Cocke, Tracking nuclear wave-packet dynamics in molecular oxygen ions with few-cycle infrared laser pulses, Phys. Rev. A 82, 013408 (2010).
@@
- Erattupuzha et al. [2016] S. Erattupuzha, S. Larimian, A. Baltuška, X. Xie, and M. Kitzler, Two-pulse control over double ionization pathways in CO2, The Journal of Chemical Physics 144, 024306 (2016).
- Rudenko et al. [2016] A. Rudenko, V. Makhija, A. Vajdi, T. Ergler, M. Schürholz, R. K. Kushawaha, J. Ullrich, R. Moshammer, and V. Kumarappan, Strong-field-induced wave packet dynamics in carbon dioxide molecule, Faraday Discussions 194, 463 (2016).
@@
- Lévêque et al. [2013] C. Lévêque, A. Komainda, R. Taïeb, and H. Köppel, Ab initio quantum study of the photodynamics and absorption spectrum for the coupled and states of SO2, The Journal of Chemical Physics 138, 044320 (2013).
@@
- Chen et al. [2023] L. Chen, E. Wang, W. Zhao, M. Gong, X. Shan, and X. Chen, Fragmentation of SO () induced by 1 keV electron collision, The Journal of Chemical Physics 158, 054301 (2023).
- Lin et al. [2020] K. Lin, X. Hu, S. Pan, F. Chen, Q. Ji, W. Zhang, H. Li, J. Qiang, F. Sun, X. Gong, H. Li, P. Lu, J. Wang, Y. Wu, and J. Wu, Femtosecond resolving photodissociation dynamics of the SO2 molecule, The Journal of Physical Chemistry Letters 11, 3129 (2020).
@@
- Rösch et al. [2023] D. Rösch, Y. Xu, H. Guo, X. Hu, and D. L. Osborn, SO2 photodissociation at 193 nm directly forms S(3P) + O2(): Implications for the archean atmosphere on earth, The Journal of Physical Chemistry Letters 14, 3084 (2023).
@@
- Goll et al. [2006] E. Goll, G. Wunner, and A. Saenz, Formation of ground-state vibrational wave packets in intense ultrashort laser pulses, Phys. Rev. Lett. 97, 103003 (2006).
- Fang and Gibson [2008] L. Fang and G. N. Gibson, Strong-Field Induced Vibrational Coherence in the Ground Electronic State of Hot , Phys. Rev. Lett. 100, 103003 (2008).
- Wei et al. [2017] Z. Wei, J. Li, L. Wang, S. T. See, M. H. Jhon, Y. Zhang, F. Shi, M. Yang, and Z.-H. Loh, Elucidating the origins of multimode vibrational coherences of polyatomic molecules induced by intense laser fields, Nat. Commun. 8, 735 (2017).
- Matsuda et al. [2014] A. Matsuda, E. J. Takahashi, and A. Hishikawa, Time-resolved laser Coulomb explosion imaging using few-cycle intense laser pulses: Application to exploding in highly charged states, Journal of Electron Spectroscopy and Related Phenomena 195, 327 (2014).
- Makhija et al. [2016] V. Makhija, X. Ren, D. Gockel, A.-T. Le, and V. Kumarappan, Orientation Resolution through Rotational Coherence Spectroscopy (2016), arXiv:1611.06476 [physics.atom-ph] .
- Wangjam et al. [2021] T. N. Wangjam, H. V. S. Lam, and V. Kumarappan, Strong-field ionization of the triplet ground state of O2, Phys. Rev. A 104, 043112 (2021).
@@
- Magrakvelidze et al. [2014] M. Magrakvelidze, A. Kramer, K. Bartschat, and U. Thumm, Complementary imaging of the nuclear dynamics in laser-excited diatomic molecular ions in the time and frequency domains, Journal of Physics B: Atomic, Molecular and Optical Physics 47, 124003 (2014).
@@
- Jahnke et al. [2025] T. Jahnke, S. Mai, S. Bhattacharyya, K. Chen, R. Boll, M. E. Castellani, S. Dold, U. Frühling, A. E. Green, M. Ilchen, R. Ingle, G. Kastirke, H. V. S. Lam, F. Lever, D. Mayer, T. Mazza, T. Mullins, Y. Ovcharenko, B. Senfftleben, F. Trinter, Atia-Tul-Noor, S. Usenko, A. S. Venkatachalam, A. Rudenko, D. Rolles, M. Meyer, H. Ibrahim, and M. Gühr, Direct observation of ultrafast symmetry reduction during internal conversion of 2-thiouracil using coulomb explosion imaging, Nature Communications 16, 2074 (2025).